# BiSSL: Enhancing the Alignment Between Self-Supervised Pretraining and Downstream Fine-Tuning via Bilevel Optimization

### v4

@@ A.2 Derivation of the Implicit Jacobian
  • $G$ is convex such that $\nabla_{{\bf\it\theta}_{P}}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))=\mathbf{0}$ is fulfilled for every ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$.
  • The Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ exists and is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$.
Then, the IJ $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be explicitly expressed by (16). The authors acknowledge that an explicit expression for the IJ without fixing ${\bf\it\phi}_{P}$ is achievable, though this is left for future exploration.
A.3 Distinction from Bilevel Optimization in Meta-Learning
While bilevel optimization (BLO) has been applied in meta-learning frameworks such as MAML [31], Sign-MAML [72] and iMAML [30], BiSSL represents a distinct application and implementation of BLO, tailored for the challenges of self-supervised learning (SSL). In the aforementioned works, BLO is primarily utilized to address few-shot learning scenarios, focusing on efficiently adapting models to new tasks with minimal labeled data. Conversely, BiSSL applies BLO to concurrently manage the more complex task of self-supervised pretext pretraining on unlabeled with downstream fine-tuning on labeled data. Another key distinction is that in meta-learning, the upper- and lower-level objectives are closely related, with the upper-level objective formulated as a summation of the lower-level tasks. In contrast, BiSSL involves fundamentally distinct objectives at each level, utilizing separate datasets and tasks for pretraining and fine-tuning. This design allows BiSSL to better align the pretrained model with the requirements of a specific downstream task. Conversely, the BLO in meta-learning aims to broadly generalize across a wide range of tasks, prioritizing adaptability rather than task-specific optimization. Lastly, unlike BiSSL, the meta-learning frameworks reinitialize the lower-level backbone parameters with a copy of the upper-level parameters at every iteration.
Appendix B Default Experiment Details
@@ B.1 Dataset Partitions
The Caltech-101 [61] dataset does not come with a pre-defined train/test split, so the same convention as previous works is followed [1, 73, 74], where 30 random images per class are selected for the training partition, and the remaining images are assigned for the test partition. For the DTD [59] and SUN397 [55] datasets, which offer multiple proposed train/test partitions, the first splits are used, consistent with the approach in [1].
For downstream hyperparameter optimization, portions of the training partitions from each respective labeled dataset are designated as validation datasets. The FGVC Aircraft [57], Oxford 102 Flowers [62], DTD, and Pascal VOC 2007 [58] datasets already have designated validation partitions. For all the remaining labeled datasets, the validation data partitions are randomly sampled while ensuring that class proportions are maintained. Roughly $20\%$ of the training data is allocated for validation.
@@ B.2 Downstream Task Fine-Tuning of the Baseline Setup
In Table 3, the learning rates and weight decays used for each respective downstream dataset of the experiments described in Section 4.2.1 are outlined.
@@ B.3 Downstream Head Warmup and Upper-level of BiSSL
The first term of the upper-level gradient (7) is approximated using the Conjugate Gradient (CG) method [47, 48]. Our implementation follows a similar structure to that used in [30], employing $N_{c}=5$ iterations and a dampening term $\lambda_{\text{damp}}=10$. Given matrix $A$ and vector $\mathbf{v}$, the CG method iteratively approximates $A^{-1}\mathbf{v}$, which requires evaluation of multiple matrix-vector products $A\mathbf{d}_{1}$, $\ldots$, $A\mathbf{d}_{N_{c}}$. In practice, storing the matrix $A$ (in our case, the Hessian $\nabla^{2}_{{\bf\it\theta}_{P}}\mathcal{L}^{P}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})$) in its full form is often infeasible. Instead, a function that efficiently computes the required matrix-vector products instead of explicitly storing the matrix is typically utilized. For transparency, the function employed in our setup is detailed in Algorithm 2. This approach ensures that the output of the CG algorithm is an approximation of the inverse Hessian-vector product in the first term of Equation (7) as intended.
@@
1:Input: Input vector $\mathbf{v}$. Model parameters ${\bf\it\theta}_{P}$, ${\bf\it\phi}_{P}$. Training objective $\mathcal{L}^{P}$. Lower-level data batche $\mathbf{x}$. Regularization weight $\lambda$ and dampening $\lambda_{\text{damp}}$.
@@
3:$\pi({\bf\it\theta}_{P})\leftarrow\quantity(\nabla_{\boldsymbol{\theta}}\mathcal{L}^{P}\left(\boldsymbol{\theta},\boldsymbol{\phi}_{P};\mathbf{x}\right)|_{\boldsymbol{\theta}=\boldsymbol{\theta}_{P}})^{T}\mathbf{v}$
4:$\mathbf{g}\leftarrow\nabla_{{\bf\it\theta}}\pi({\bf\it\theta})\big{|}_{\boldsymbol{\theta}=\boldsymbol{\theta}_{P}}$$\triangleright$ Memory efficient calculation of $\nabla_{{\bf\it\theta}}^{2}\mathcal{L}^{P}({\bf\it\theta},{\bf\it\phi}_{P};\mathbf{x})|_{{\bf\it\theta}={\bf\it\theta}_{P}}\mathbf{v}$.
5:$\mathbf{y}\leftarrow\mathbf{v}+\frac{1}{\lambda+\lambda_{\text{damp}}}\mathbf{y}$
@@ B.5 Downstream Fine-Tuning after BiSSL
Food $0.0037$ $0.00018$ $0.0496$ $0.00096$
CIFAR10 $0.001$ $0.00278$ $0.0018$ $0.00014$
CIFAR100 $0.0019$ $0.00002$ $0.0008$ $0.0013$
CUB200 $0.0005$ $0.00331$ $0.0007$ $0.00947$
SUN397 $0.0015$ $0.00001$ $0.0009$ $0.00062$
Cars $0.0289$ $0.0075$ $0.0316$ $0.00632$
Aircrafts $0.0002$ $0.0059$ $0.0003$ $0.00001$
VOC07 $0.0002$ $0.00019$ $0.0004$ $0.0003$
DTD $0.0003$ $0.00002$ $0.0013$ $0.0021$
Pets $0.0014$ $0.00005$ $0.0002$ $0.00039$
Caltech101 $0.0016$ $0.00042$ $0.0005$ $0.00022$
Flowers $0.0189$ $0.00143$ $0.0057$ $0.00207$
B.6 Object Detection
@@
The experimental setup primarily follows the main implementation specified in Section 4.2, with minor modifications specified here.
@@
We utilize a Faster R-CNN with a ResNet-50 C4 backbone [67] as the downstream model. The downstream data augmentations involve rescaling the images so that their longest edge is between196196196196 and $320$ pixels, followed by normalization. Downstream fine-tuning is conducted for $50$ epochs with a batch size of $16$. The random hyperparameter grid search is performed across $50$ distinct configurations of learning rates and weight decays, sampled log-uniformly within the intervals $10^{-4}$ to $10^{-1}$ and $10^{-6}$ to $10^{-2}$, respectively.
Important for the context of BiSSL, we extend the pretext head parameters ${\bf\it\phi}_{P}$ to further include the latter layers of the ResNet-50 that are not part of the ResNet-50 C4 backbone. Lastly, the linear head warm-up is conducted for 5 epochs.

### v5

@@ A.1 Derivation of the Implicit Jacobian
  • The regularization objective $r$ is strongly convex.
  • The objective $\mathcal{L}^{P}$ is twice-differentiable and $L$-smooth, with $\lambda$ accordingly chosen so that $G$ is convex.
  • The Hessian matrix $\nabla_{{\bf\it\theta}_{P}}^{2}G({\bf\it\theta}_{D},{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}))$ is invertible for all ${\bf\it\theta}_{D}\in\mathbb{R}^{L}$.
Then, the IJ $\frac{\text{d}{\bf\it\theta}_{P}^{*}({\bf\it\theta}_{D})}{\text{d}{\bf\it\theta}_{D}}^{T}$ can be explicitly expressed by (9). The authors further acknowledge that an explicit expression for the IJ without fixing ${\bf\it\phi}_{P}$ is achievable, though this is left for future exploration. The assumptions in the final two bullets above may not strictly hold in practice with deep neural networks, yet the resulting gradient expressions have nonetheless proven practically beneficial both in this and prior work.
A.2 Distinction from Bilevel Optimization in Meta-Learning
While bilevel optimization (BLO) has been applied in meta-learning frameworks such as MAML (Finn et al., 2017), Sign-MAML (Fan et al., 2021) and iMAML (Rajeswaran et al., 2019), BiSSL represents a distinct application and implementation of BLO, tailored for the challenges of self-supervised learning (SSL). In the aforementioned works, BLO is primarily utilized to address few-shot learning scenarios, focusing on efficiently adapting models to new tasks with minimal labeled data. Conversely, BiSSL applies BLO to concurrently manage the more complex task of self-supervised pretext pretraining on unlabeled data with downstream fine-tuning on labeled data. Another key distinction is that in meta-learning, the upper- and lower-level objectives are closely related, with the upper-level objective formulated as a summation of the lower-level tasks. In contrast, BiSSL involves fundamentally distinct objectives at each level, utilizing separate datasets and tasks for pretraining and fine-tuning. This design allows BiSSL to better align the pretrained model with the requirements of a specific downstream task. Conversely, the BLO in meta-learning aims to broadly generalize across a wide range of tasks, prioritizing adaptability rather than task-specific optimization.
Appendix B Experimental Details
@@ B.1 Dataset Partitions
The Caltech-101 (Li et al., 2022a) dataset does not come with a pre-defined train/test split, so the same convention as previous works is followed (Chen et al., 2020; Donahue et al., 2014; Simonyan and Zisserman, 2014), where 30 random images per class are selected for the training partition, and the remaining images are assigned for the test partition. For the DTD (Cimpoi et al., 2014) and SUN397 (Xiao et al., 2010) datasets, which offer multiple proposed train/test partitions, the first splits are used, consistent with the approach in (Chen et al., 2020).
For downstream hyperparameter optimization, portions of the training partitions from each respective labeled dataset are designated as validation datasets. The FGVC Aircraft (Maji et al., 2013), Oxford 102 Flowers (Nilsback and Zisserman, 2008), DTD, and Pascal VOC 2007 (Everingham et al., 2010) datasets already have designated validation partitions. For the remaining labeled datasets, we reserve roughly $20\%$ of the training data for validation. For CityScapes (Cordts et al., 2016), this split is applied separately within each city, while for the other datasets, validation partitions are randomly sampled while ensuring that the class proportions are preserved.
@@ B.2 Downstream Task Fine-Tuning of the Baseline Setup
A random grid search of $100$ hyper-parameter configurations for the base learning rates and weight decays is conducted, where one model is fine-tuned for each configuration. Base learning rates and weight decays are log-uniformly sampled over the ranges of $10^{-4}$ to 11 and $10^{-5}$ to $10^{-2}$, respectively. Validation data accuracy is evaluated after each epoch. The hyper-parameter configuration yielding the best balance between high validation accuracy and low validation loss is considered the optimal hyper-parameter configuration. [^2] In certain scenarios during the experiments, the configuration that achieved the highest validation accuracy also yielded a notably higher relative validation loss. To ensure better generalizability, an alternative configuration with a more favorable trade-off was selected in these cases. The corresponding optimal hyper-parameters for each downstream dataset are documented in Table 5. For subsequent evaluation on the test data, we train $10$ models with different random seeds, each using the considered optimal hyper-parameter configurations. During the training of each respective model, the model parameters are stored after each epoch if the top-1 validation accuracy (or 11-point mAP for the VOC07 dataset) has increased compared to the previous highest top-1 validation accuracy achieved during training. Top-1 and top-5 test data accuracies (or 11-point mAP for the VOC07 dataset) are evaluated for each of the $10$ models, from which the calculated means and standard deviations of these accuracies are documented.
@@ B.3 Downstream Head Warmup and Upper-level of BiSSL
The first term of the upper-level gradient (7) is approximated using the Conjugate Gradient (CG) method (Nazareth, 2009; Shewchuk, 1994). Our implementation follows a similar structure to that used in Rajeswaran et al. (2019), employing $N_{c}=5$ iterations and a dampening term $\lambda_{\text{damp}}=10$. Given matrix $A$ and vector $\mathbf{v}$, the CG method iteratively approximates $A^{-1}\mathbf{v}$, which requires evaluation of multiple matrix-vector products $A\mathbf{d}_{1}$, $\ldots$, $A\mathbf{d}_{N_{c}}$. In practice, storing the matrix $A$ (in our case, the Hessian $\nabla^{2}_{{\bf\it\theta}_{P}}\mathcal{L}^{P}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\phi}_{P})$) in its full form is often infeasible. Instead, a function that efficiently computes the required matrix-vector products instead of explicitly storing the matrix is typically utilized. For transparency, the function employed in our setup is detailed in Algorithm 2. This approach ensures that the output of the CG algorithm is an approximation of the inverse Hessian-vector product in the first term of Equation (7) as intended.
@@
1:Input: Input vector $\mathbf{v}$. Model parameters ${\bf\it\theta}_{P}$, ${\bf\it\phi}_{P}$. Training objective $\mathcal{L}^{P}$. Lower-level data batch $\mathbf{x}$. Regularization weight $\lambda$ and dampening $\lambda_{\text{damp}}$.
@@
[…]
@@ B.5 Downstream Fine-Tuning after BiSSL
Food $0.00059$ $0.00008$ $0.00018$ $0.00019$
CIFAR10 $0.00097$ $0.00258$ $0.00098$ $0.00039$
CIFAR100 $0.00252$ $0.00046$ $0.00042$ $0.00292$
CUB200 $0.00043$ $0.00251$ $0.00066$ $0.00075$
SUN397 $0.00079$ $0.00001$ $0.00063$ $0.00029$
Cars $0.05706$ $0.00412$ $0.06221$ $0.00317$
Aircrafts $0.00015$ $0.00003$ $0.00025$ $0.00056$
VOC07 $0.00022$ $0.00592$ $0.00029$ $0.004$
DTD $0.00062$ $0.00002$ $0.00028$ $0.00008$
Pets $0.00153$ $0.00314$ $0.00042$ $0.00002$
Caltech101 $0.00079$ $0.00005$ $0.00194$ $0.00003$
Flowers $0.01768$ $0.0009$ $0.00058$ $0.00006$
B.6 Statistical Significance Tests
As described in Section B.2, each documented accuracy is obtained by fine-tuning $10$ models with different random seeds. When comparing two sets of observations, such as FT and BiSSL+FT in Table 1, this yields two groups of $10$ values each, for a total of $20$ observations. Under the standard assumption that the values within each group are independent and identically distributed, we use a two-sample permutation test to assess whether the average accuracy of each group differs significantly.
Under the null hypothesis that all $20$ values come from the same distribution, the test pools the values, randomly assigns $10$ to each group, and computes the resulting difference between group averages for each permutation. Since there are $\binom{20}{10}=184,756$ unique permutations, we perform the exact test by evaluating all possible assignments.
The resulting p-value is the proportion of permutations where the difference of averages is at least as large as the observed difference. We consider p-values below $0.01$ to indicate a statistically significant difference in average accuracy between the two groups.
B.7 Object Detection and Semantic Segmentation
@@
The experimental setup primarily follows the main implementation specified in Section 4.2, with the specific modifications made specified here.
Object Detection
@@
We utilize a Faster R-CNN with a ResNet-50 C4 backbone (Ren et al., 2015) as the downstream model architecture. In line with Grill et al. (2020); He et al. (2020); Bardes et al. (2022), we assign the box prediction head as the conv5 stage of the ResNet-50 followed by global pooling. We use anchors of size $32$, $64$, $128$ and $256$ with aspect ratios $0.5$, $1.0$ and $2.0$. The downstream data augmentations involve rescaling the images so that their longest edge is between $196$ and $320$ pixels, followed by normalization. Downstream fine-tuning is conducted for $50$ epochs with a batch size of $16$. The random hyperparameter grid search is performed across $50$ distinct configurations of learning rates and weight decays, sampled log-uniformly within the intervals $10^{-4}$ to $10^{-1}$ and $10^{-6}$ to $10^{-2}$, respectively. The linear head warm-up, conducted prior to BiSSL, is executed for 5 epochs.
We note that, since the downstream head is inserted between intermediate stages of the backbone, this setup does not fully comply with the notation in Section 3.1, where the downstream head is assumed to operate solely on the backbone’s final output layer. Nonetheless, the derived gradient expression (7) retains its validity in this setting, as it does not depend on any part of the downstream head but only requires a one-to-one correspondence between the parameter-wise components between the levels. In particular, the conv5 stage parameters of the lower-level backbone can still be directly substituted into the corresponding conv5 stage of the upper-level backbone.
Semantic Segmentation
For semantic segmentation tasks, we adopt an FCN-based architecture (Long et al., 2015a) similar to the setup used in Grill et al. (2020); He et al. (2020). We use the ResNet-50 backbone output prior to global pooling, and the $3\crossproduct 3$ convolutions in the conv5 block are reassigned dilation $2$ and stride $1$. The downstream head is composed of two $3\crossproduct 3$ convolutions, each with dilation $6$ and $256$ channels, with batch normalization and ReLU activation applied in between. A final $1\crossproduct 1$ convolution performs the per-pixel classification required for segmentation, resulting in a total stride of $16$ (corresponding to the FCN-16 setup (Long et al., 2015a)).
Data augmentation includes resizing such that the shorter image side is $192$ pixels, followed by random horizontal flipping with probability $0.5$ and normalization. The fine-tuning hyperparameter search is identical to the object detection setup, and the linear head warm-up is similarly conducted for $5$ epochs. Although some structural modifications are applied to the backbone, similar to the object detection configuration, these changes do not introduce any compatibility issues with BiSSL.
B.8 Masked Autoencoder with ViT Backbone
Due to computational constraints, we adopt a relatively lightweight configuration with a ViT-S backbone (Touvron et al., 2021; Heo et al., 2021) and an $8\crossproduct 8$ patch size. The decoder matches the original MAE implementation (He et al., 2022). Pretext pretraining was run for $500$ epochs using the AdamW (Loshchilov and Hutter, 2019) optimizer with a cosine-decayed learning rate starting at $0.0005$, weight decay of $0.05$, and momentum coefficients of $\beta_{1}=0.9$ and $\beta_{2}=0.95$. The lower-level BiSSL stage used the same optimizer but with an initial learning rate of $0.0001$. For both upper-level optimization and downstream fine-tuning, we applied a $10\%$ drop path rate and layer-wise learning rate decay with a factor of $0.75$. We conducted the random hyperparameter grid search over $50$ different configurations. All other experimental settings match those described in Section 4.2.2.
B.9 Additional Baseline Comparisons
This section outlines the complete implementation details for each newly added baseline according the experiments summarized in Table 3 of Section 4.4.1. In all experiments within this section involving self-supervised pretraining, we use SimCLR as the pretext task.
B.9.1 Random Initialization, Linear Probing and Zero-Initialized Classification Head
The first baseline uses randomly initialized backbone weights instead of the SimCLR-pretrained backbone. The linear probing baseline freezes the backbone weights, and the zero-initialized head baseline sets the classification head parameters to zero rather than random values. As prior work suggests that randomly initialized heads may cause a distortion of pretrained features during fine-tuning (Kumar et al., 2022; Yang et al., 2022), we hypothesize that zero-initializing the downstream head may mitigate such loss. For each result, the hyperparameter grid search is conducted over $25$ different combinations. All other aspects of the experimental setup follow Section 4.2.1. Results are included in the respective "Random Initialization, "Linear Probing" and "FT (Zero Head Init)" rows of Table 3.
B.9.2 NoisyTune
NoisyTune (Wu et al., 2022) applies layer-wise Gaussian noise to the backbone weights prior to fine-tuning, with the noise magnitude determined by each layer’s statistics. We use a noise scale of $0.015$, and conduct the hyperparameter grid search over $50$ different combinations. All other aspects of fine-tuning follow Section 4.2.1. Results are reported in the "NoisyTune" row of Table 3.
B.9.3 Continued Pre-Training on Downstream and Pretext Dataset Mixture
For each downstream dataset, we resumed self-supervised pretraining after adding the downstream data to the original pretraining data pool. This baseline thus continues pretraining on a mixture of pretext and downstream data before fine-tuning.
Given the current BiSSL setting conducts what is roughly equivalent to $8$ pretext epochs, we here continued SSL pretraining for $10$ epochs using a cosine-decayed learning rate starting at $1.0$. All other pretraining settings match those described in Section 4.2.1. The subsequent fine-tuning follows the standard setup, but with a hyperparameter grid search over $25$ combinations. Results for this baseline are reported in the “Pretraining on Data Mix + FT” row of Table 3.
Accuracy $w=\mathbf{0}$ $w=\mathbf{0.05}$ $w=\mathbf{0.1}$ $w=\mathbf{0.25}$ $w=\mathbf{0.5}$ $w=\mathbf{0.75}$ FT BiSSL+FT
Top-1 $60.1\pm 0.3$ $61.6\pm 0.2$ $61.7\pm 0.2$ $\underline{62.2\pm 0.3}$ $60.8\pm 0.3$ $60.3\pm 0.5$ $60.3\pm 0.9$ $\mathbf{64.2\pm 0.4}$
Top-5 $\underline{86.3\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.2\pm 0.3}$ $\underline{86.1\pm 0.3}$ $\underline{86.2\pm 0.3}$ $85.7\pm 0.3$ $85.8\pm 0.6$ $\mathbf{87.8\pm 0.3}$
B.9.4 Weighted Sum of Pretext and Downstream Objectives
This baseline is trained by solving the single-level optimization problem:
We use the lower-level optimizer configuration from BiSSL (see Section 4.2.2) and train for $4000$ steps, matching the total number of upper-level iterations in the default BiSSL configuration. Fine-tuning follows the main setup but with a grid search over 50 hyperparameter combinations. Backbone gradients for the two loss terms are computed separately on their respective distinct mini-batches, scaled by 1w1-w and $w$ respectively, and then summed prior to the update step.
To determine the suitable size of $w$, we conducted a sweep on the DTD dataset over different values of $w$ in the range between 0 and $1$, evaluating both top-1 and top-5 classification performance. For each value of $w$, we conduct the subsequent fine-tuning hyperparameter grid search over $25$ different combinations, otherwise the fine-tuning setup remains identical to as described in Section 4.2.1. The results of the sweep are shown in Table 8.
The results imply that $w=0.25$ provides the best top-1 classification accuracy. In contrast to BiSSL, the top-5 accuracy does generally not show a significant increase compared to conventional fine-tuning, regardless of the value of $w$. Setting $w=0$ essentially reduces the training to pretext-only learning, reflected by comparable performance to conventional fine-tuning (the "FT" column), with slightly lower top-1 accuracy and slightly higher top-5 accuracy. Based on the results, we consider $w=0.25$ as the optimal value for this baseline.
Using $w=0.25$, we summarize the baseline experiments in the "Pretext-Downstream Sum + FT" row in Table 3.
B.9.5 Discarded First Term of Upper-Level Objective in BiSSL
For this baseline, we discard the first term of the upper-level objective in (2), which contains the IJ. This blocks the propagation of lower-level gradient information to the upper-level, hence instead resembling a more conventional alternating optimization scheme where only parameter values are shared. The rest of the setup follows the default configuration in Section 4.2.2, except that the fine-tuning hyperparameter grid search was conducted for $50$ different combinations. The results are provided in the "BiSSL (IJ Discarded) + FT"-row of Table 3.
B.10 Backbone Alignment Analysis
B.10.1 Cosine Similarities
Recall the notation in Section 3.1, where we let $\bf\it\theta$} denote the backbone parameters obtained after pretext task training, and ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ the backbone after applying BiSSL. We now further define ${\bf\it\theta}^{FT}_{i}$ and ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})^{FT}_{i}$ for $i=1,\dots,10$ as the fine-tuned backbone parameters of the pretext and BiSSL models, respectively, which we achieve $10$ of for each downstream dataset (recall Section B.2). Lastly, let
$cs_{N}(\mathbf{x},\mathbf{y})=\frac{\mathbf{x}^{T}\mathbf{y}}{\norm{\mathbf{x}}_{2}^{2}\norm{\mathbf{y}}_{2}^{2}},\quad\mathbf{x},\mathbf{y}\in\mathbb{R}^{N}/\{\mathbf{0}\}$ (11)
denote the cosine similarity of $N$-dimensional vectors.
Table 9: Parameter-wise and feature-wise cosine similarities (multiplied by $100$ for better readability) between SimCLR-pretrained and BiSSL-trained backbones with their respective fine-tuned counterparts for each dataset. In every case, the BiSSL/Fine-Tuned cosine similarities are statistically significantly higher, as indicated by bold font.
Dataset Cosine Similarity (Parameters) Cosine Similarity (Features)
Pretext/Fine-Tuned BiSSL/Fine-Tuned Pretext/Fine-Tuned BiSSL/Fine-Tuned
Food $15.88\pm 0.22$ $\mathbf{99.79\pm 0.03}$ $16.77\pm 0.43$ $\mathbf{93.50\pm 0.37}$
[…]
CUB200 $17.01\pm 0.20$ $\mathbf{99.99\pm 0.00}$ $26.74\pm 0.26$ $\mathbf{96.28\pm 0.02}$
[…]
Cars $24.42\pm 0.11$ $\mathbf{26.64\pm 0.10}$ $28.78\pm 0.34$ $\mathbf{65.01\pm 0.14}$
Aircrafts $65.74\pm 0.15$ $\mathbf{99.95\pm 0.01}$ $31.42\pm 0.27$ $\mathbf{90.97\pm 0.24}$
[…]
Average (Fig. 3) $76.46$ $\mathbf{93.82}$ $43.80$ $\mathbf{90.63}$
The purpose of these evaluations is then to measure how aligned the pretext pretrained backbones and their fine-tuned counterparts ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ and ${\bf\it\theta}^{FT}_{i}$ are compared to the BiSSL-trained backbones and their fine-tuned counterparts ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ and ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})^{FT}_{i}$. Specifically, we first compute the parameter-wise cosine similarities
$cs_{L}({\bf\it\theta},{\bf\it\theta}^{FT}_{i})\quad\quad\text{and}\quad\quad cs_{L}({\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D}),{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})^{FT}_{i}),$ (12)
for $i=1,\ldots,10$. For each downstream dataset, the mean and standard deviation of these values are reported in the left column of Table 9.
We also compute the average cosine similarity between output features of the respective backbones for each downstream dataset on its test partition $\mathcal{D}^{D_{\text{test}}}=\{\mathbf{x}_{k},\mathbf{y}_{k}\}_{k=0}^{|\mathcal{D}^{D_{\text{test}}}|}$, given by
$\frac{1}{|\mathcal{D}^{D_{\text{test}}}|}\sum_{\mathbf{x},\mathbf{y}\in\mathcal{D}^{D_{\text{test}}}}cs_{M}(f_{{\bf\it\theta}}(\mathbf{x}),f_{{\bf\it\theta}^{FT}_{i}}(\mathbf{x}))\quad\quad\text{and}\quad\quad\frac{1}{|\mathcal{D}^{D_{\text{test}}}|}\sum_{\mathbf{x},\mathbf{y}\in\mathcal{D}^{D_{\text{test}}}}cs_{M}(f_{{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})}(\mathbf{x}),f_{{\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})^{FT}_{i}}(\mathbf{x})),$ (13)
for $i=1,\ldots,10$. During the evaluation, it is important to note that the batch normalization layers (Ioffe and Szegedy, 2015) of the backbones utilize the running means and variances inferred during training. Since the pretext and BiSSL backbones $\bf\it\theta$}, ${\bf\it\theta}_{P}^{\ast}({\bf\it\theta}_{D})$ have not been exposed to the downstream datasets during training, their batch normalization statistics may likely be suboptimal for the downstream datasets. To address this, the downstream training dataset is divided into batches of $256$ samples, and roughly $100$ batches are then forward-passed through these backbones, prior to calculating the output features in (13). This procedure updates the BN statistics to better reflect the downstream data, enabling a fairer comparison of the learned representations. The resulting feature-wise cosine similarities for each downstream dataset are reported in the right column of Table 9.
B.10.2 Visual inspection of Latent Features
In same manner as in Section B.10.1, we first calibrate the batch normalization running statistics by forward-passing roughly $100$ downstream training batches through the backbones. We then extract the backbone output features for all samples in the test partition, and reduce their dimensionality to two using t-Distributed Stochastic Neighbor Embedding (t-SNE) (Cieslak et al., 2020). Additional plots on other downstream datasets are provided in Section C.7.